### v1

# Exploring the extent of similarities in software failures across industries using LLMs

Abstract
Background/Motivation:
The rapid evolution of software development necessitates enhanced safety measures. Extracting information about software failures from companies is becoming increasingly more available through news articles.
Problem Statement:
This research utilizes the Failure Analysis Investigation with LLMs (FAIL) model to extract industry-specific information. Although the FAIL model’s database is rich in information, it could benefit from further categorization and industry-specific insights to further assist software engineers.
Methods:
In previous work news articles were collected from reputable sources and categorized by incidents inside a database. Prompt engineering and Large Language Models (LLMs) were then applied to extract relevant information regarding the software failure. This research extends these methods by categorizing articles into specific domains and types of software failures. The results are visually represented through graphs.
Results:
The analysis shows that throughout the database some software failures occur significantly more often in specific industries. This categorization provides a valuable resource for software engineers and companies to identify and address common failures.
Conclusion:
This research highlights the synergy between software engineering and Large Language Models (LLMs) to automate and enhance the analysis of software failures. By transforming data from the database into an industry specific model, we provide a valuable resource that can be used to identify common vulnerabilities, predict potential risks, and implement proactive measures for preventing software failures. Leveraging the power of the current FAIL database and data visualization, we aim to provide an avenue for safer and more secure software in the future.
Keywords: Software Failure, Prompt Engineering
Introduction
The software industry’s dynamic nature demands continual advancements in safety protocols and failure analysis techniques. Current methodologies for studying software failures often face set backs due to restrictive data access and privacy concerns. The news seems to be the optimal route for collecting information using Large Language Models (LLMs).
The FAIL database provides a well established foundation, compiling detailed reports of software failures from various news outlets. However, there exists a significant opportunity to delve deeper into the contextual specifics of these failures, such as the industries most affected and the predominant types of failures in each industry.
In response to these needs, this study expanded the use of the FAIL model by categorizing failures not just broadly but within specific industry contexts. This involved enhancing data extraction techniques through prompt engineering with LLMs, aiming to achieve a more nuanced understanding of where and why these failures occur.
Our analysis demonstrated that some types of software failures are notably more prevalent in certain industries. By identifying patterns, we provide insights that can drive targeted improvements in software practices and policies, potentially leading to substantial reductions in software failures.
Background And Significance
The Securities and Exchange Commission (SEC) recently made a ruling requiring public companies to disclose cyber-security related failures. This means that any breach in security deemed to have potential significant impact or harm must be reported within 4 business days of the incident. Public companies must also create reports annually of cyber-security risk management strategies, and governance [7]. With this increased transparency requirement in the industry, the FAIL model, and industry specific insights will only become more useful and prevalent in the future.
Related Work
This work is directly correlated to the work done by FAIL: Analyzing Software Failures from the News Using LLMs [1], materializing a way to automate the process of gathering information on software failures from news articles. In this paper they explored an efficient approach to automate the process of collecting information on recent software failures from the news. They collected information only from articles that the LLMs deemed to be having sufficient enough information to analyze. They then categorized this data into summaries, postmortems, and taxonomy. They collected valuable information, and the information is informative. The information could benefit from further insights into industry-specific failures.
Methodology
We began by creating a prompt from which we collect data from the database, and categorize the data into industry-specific failures. The initial taxonomy was adopted from [8, 9, 10] which give taxonomies for categorizing software failures. The temperature of the LLM that we used was manually set to 0 to increase consistency, and we used chatgpt-3.5-turbo. Initially the prompt contained general instructions, such as ”categorize these prompts by type of software failure.” However the prompt’s performance was sub optimal. Effective prompts need to be clear and concise to guide the Large Language Model (LLM) in producing the desired output [2]. The initial prompt generated results with high variability or ”randomness”. To improve accuracy, we revised the prompt with a specific list of common software failures manually curated based on descriptions in the database, and taxonomies widely used in the industry for categorizing software failures. [10, 9] We ensured the prompt only returned the type of software failure. This adjustment yielded more consistent results by the LLM, but our approach was still producing some incorrect categorizations. After refining the prompt and using few-shot prompting, the prompt performed significantly better. Through the use of a confusion matrix and a small dataset of 90 articles, the overall average accuracy of the prompt was 76 percent (reference Figure 1 for an example of the matrix used). With this accuracy we let the script run through the database, and categorize each software failure by industry. Results are shown visually through bar graphs.
Figure 1: Confusion matrix used for analysis.
Results
Overview
This section presents the analysis of software failure frequencies across different industries.
Finance Industry
The graph in figure 2(a) shows that security-related failures are the most common in the finance industry, with Security vulnerabilities being significantly more frequent than other types of failures.
Healthcare Industry
As shown in figure 2(b) shows security vulnerabilities are the most common in the healthcare industry. This could be due to the healthcare industry storing a vast amount of patient data with inadequate security.
Information Industry
As shown in figure 2(c) shows security vulnerabilities are most common in the information industry. The vast amount of data within the IT (information Technology) industry increases the risk of security related failures.
Knowledge Industry
As shown in figure 2(d) shows security vulnerabilities are most common in the Knowledge industry. This industry includes educational institutions, research organizations, and knowledge-based enterprises. Security related failures could be happening more frequently due to the openness of sharing information in an educational environment.
Transportation Industry
As shown in figure 2(e) shows functionality bugs are most common in the Transportation industry. This industry relies heavily on complex logistics and random variables, which could in effect be the reason we are seeing more functionality bugs.
Entertainment Industry
As shown in figure 2(f) shows security vulnerabilities are most common in the Entertainment industry. This industry has a large audience, and often uses third party software to manage data, which could lead to increased vulnerabilities.
Government Sector
As shown in figure 3(a) shows security vulnerabilities are most common in the Government Sector. With the government sector handling a large range of sensitive information, it’s no surprise that security vulnerabilities are at the forefront of software failures for this domain.
Security VulnerabilityFunctionality BugData BreachOutageIntegration IssueOtherPerformance IssueUI/UX BugRegression BugNon-Software Cause005050505010010010010095959595111111113535353524242424121212125555444400003333Failure TypeNumber of failuresIn Finance Industry
(a) Graph of failures within the finance industry
Security VulnerabilityFunctionality BugData BreachOutageIntegration IssueOtherPerformance IssueUI/UX BugRegression BugNon-Software Cause00202020204040404047474747202020201919191966666666555555552222001111Failure TypeNumber of failuresIn Health Industry
(b) Graph of failures within the healthcare industry
Security VulnerabilityFunctionality BugData BreachOutageIntegration IssueOtherPerformance IssueUI/UX BugRegression BugNon-Software Cause002002002002004004004004006006006006008008008008007167167167161271271271271201201201206464646433333333313131312626262618181818161616165555Failure TypeNumber of failuresIn Information Industry
(c) Graph of failures within the information industry
Security VulnerabilityFunctionality BugData BreachOutageIntegration IssueOtherPerformance IssueUI/UX BugRegression BugNon-Software Cause0010101010202020201818181812121212555533338888141414146666000000Failure TypeNumber of failuresIn Knowledge Industry
(d) Graph of failures within the knowledge industry
Security VulnerabilityFunctionality BugData BreachOutageIntegration IssueOtherPerformance IssueUI/UX BugRegression BugNon-Software Cause001001001001002002002002008080808018718718718744441616161657575757525252529999666644442222Failure TypeNumber of failuresIn Transportation Industry
(e) Graph of failures within the transportation industry
Security VulnerabilityFunctionality BugData BreachOutageIntegration IssueOtherPerformance IssueUI/UX BugRegression BugNon-Software Cause002020202040404040606060606060606033333333242424244444999966661515151544440000Failure TypeNumber of failuresIn Entertainment Industry
(f) Graph of failures within the entertainment industry
Security VulnerabilityFunctionality BugData BreachOutageIntegration IssueOtherPerformance IssueUI/UX BugRegression BugNon-Software Cause0020202020404040406060606064646464111111111818181822227777999922223333111100Failure TypeNumber of failuresIn Government Sector
(a) Graph of failures within the government sector
Discussion
Interpretation of Results:
Our results indicate that certain software failures occur significantly more frequently in specific industries. For instance, the analysis revealed that the information industry experiences a higher incidence of security-related failures, while the transportation industry often encounters functionality issues. These findings highlight the distinct challenges faced by different sectors, underscoring the need for tailored approaches to software reliability. We can also see a trend across industries, suggesting that many different industries are facing the same software development challenges.
Our study dives deeper by quantifying these differences and providing a detailed breakdown of failure types within each industry. This granularity offers new insights into the specific vulnerabilities and common issues faced by different sectors.
Implications of Findings
  • Industry specific insights: Our study provides industry-specific insights into software failures to better cater efforts towards a specific industry.
  • Efficiency: This could allow for more efficient software maintenance and improvements when a lack of resources is present.
Policy and Best Practices
Given that security vulnerabilities are the most frequent software failure across the large majority of industries, we recommend the following policy and best practices based upon studies such as [3, 4, 5]:
  1. 1.
    Regular Security Audits: Conduct regular security audits to identify and mitigate vulnerabilities in the software. This includes code reviews, penetration testing, and compliance checks. Regular security audits can aid in catching security vulnerabilities early when a fix is more plausible.
  2. 2.
    Security Training: Provide ongoing security training for software engineers and developers to stay updated with the latest security threats and mitigation techniques. Informing the group of developers about common security breaches and fixes in the past can be helpful in preventing them in the future.
  3. 3.
    Secure Coding Practices: Implement secure coding practices, such as input validation, error handling, and secure authentication mechanisms, to prevent common vulnerabilities like SQL injection and cross-site scripting (XSS). Making sure that your coding practices are consistent, and secure is one of the most practical solutions to security vulnerabilities, as we realize software development can involve large teams all working in separate sectors. Consistency can benefit debugging, and communication pertaining to security.
  4. 4.
    Incident Response Plan: Develop and maintain a comprehensive incident response plan to quickly address and mitigate the impact of security breaches. Having a response plan will mitigate damages, and ensure that the security vulnerabilities are dealt with quickly and efficiently.
  5. 5.
    Access Control: Ensure strict access control policies are in place, with the principle of least privilege, to minimize the risk of unauthorized access to sensitive data. For more information please see [4].
  6. 6.
    Regular Updates: Keep all software and dependencies updated with the latest security patches and updates to protect against known vulnerabilities.
Limitations:
  • Limited scope: Our study uses only the data found in the FAIL database, which does not capture every failure across every industry and can be seen as limited.
  • Use of Large Language Models: Our study utilizes Large Language Models to categorize software failures. Large Language models can be inaccurate, and have the possibility of giving false information known as hallucinations.
  • Using News as a Source: The FAIL database uses news articles as a source of collecting information on software failures. News can have biases, and have only bits of information.
Conclusion
In conclusion our study highlights the importance of industry-specific software failure distinctions. By identifying the most common failures in each specific sector we can inform more directive preventive measures towards software failures. These findings have significant implications for software engineers and industry best practices, paving the way for a safer software and practices in the future.
Our research provides a foundation for future research to build upon, with the potential to enhance the FAIL model with new tools, and better informational gathering techniques. Ultimately our work contributes to a deeper understanding of software failures and supports the development of targeted solutions to prevent industry-specific software failures in the future.
Acknowledgements
PI: Jamie C Davis: For his invaluable guidance, support, and encouragement throughout this research project.
NSF CPS Frontier: Cognitive Autonomy for Human CPS: for funding this project.